# Weaker Assumptions for Asymmetric Trust

### v1

@@ 1 Introduction
Asymmetric trust models, such as those proposed by Damgård et al. [12], Alpos et al. [2], and Losa et al. [19] allow the development of distributed protocols in which each participant can operate under its own trust settings. These algorithms are built on top of asymmetric quorum systems, where each process independently defines its own quorums. The traditional consistency and availability properties must be satisfied by these systems, but as was shown by Li et al. [17], they are not enough to solve fundamental problems like reliable broadcast and consensus. Therefore, the models introduce additional assumptions on the quorums. The protocols for reliable broadcast and consensus of Alpos et al. [2] require the existence of a guild, a set of processes that contains a quorum for each of its members. Li et al. [17] and Losa et al. [19] identify similar conditions. These assumptions can be very strong and restrictive, and while they are sufficient to provide solutions to the aforementioned problems, we will show that they are not necessary. This leads to asymmetric algorithms that work under weaker assumptions, making such systems more flexible and usable.
We start by showing that current assumptions for algorithms such as reliable broadcast or consensus are so restrictive that it becomes unnecessary to use asymmetric trust models altogether. That is, given an asymmetric trust assumption that satisfies such requirements, we can derive an equivalent symmetric trust assumption such that no process will be worse off by adopting it. This result is surprising, considering that Alpos et al. [2] show that asymmetric trust is more expressive than symmetric trust.
@@ 2 Related Work
Threshold and symmetric quorum systems are fundamental tools in distributed computing, commonly used to establish the safety and liveness guarantees of various protocols [22, 20, 5, 10]. They constitute a crucial component of many systems in cloud computing [16, 15] and cryptocurrencies [13, 8, 9]. One drawback is that all participants must use the same quorum system, which restricts their freedom to choose who to trust. The asymmetric model, introduced by Damgård et al. [12] and further developed by Alpos et al. [2], represents a novel approach that generalizes the symmetric paradigm. It allows participants to make trust choices of their own, reflecting social connections or other information available to them outside the protocol. Each participant can specify its own quorum system, giving it the freedom to have its own trust assumptions. Amores-Sesar et al. [4] show that algorithms relying on symmetric quorums cannot be applied directly to the asymmetric model, as certain properties do not hold if the underlying quorums are asymmetric. This drives the need to adapt existing symmetric protocols and redefine protocol properties to fit the new model.
@@
Existing work [2] introduces asymmetric versions of relevant primitives like reliable broadcast, binary consensus and common coin. Amores-Sesar et al. [4] extend this work by developing a DAG-based consensus protocol that relies on such primitives. Losa et al. [19] present another approach to model asymmetric trust. They forego the notion of fail-prone sets and strengthen the definition of quorums, by requiring each quorum to contain a quorum for each member. Li et al. [17] expand on this model and address the topic of which quorum properties are necessary or sufficient to solve core asymmetric problems such as consensus. Sheff et al. [25] propose a variant of the Paxos protocol that incorporates heterogeneous trust assumptions. Senn and Cachin [24] explore asymmetric trust in the crash-fault scenario and show that, in this setting, asymmetric trust reduces to symmetric trust.
@@
Asymmetric trust introduces novel challenges to the design of distributed protocols. One significant hurdle is that of processes that have wrong friends, that is, the existence of processes that are correct and behave according to the underlying algorithm, but whose trust assumptions do not capture the actual faulty processes. Alpos et al. [2] use the term naive to refer to such participants and point out that their presence can lead to safety and liveness issues in asymmetric protocols. To address this, they introduce the concept of a guild, a set of processes that contains a quorum for each member, and provide algorithms for reliable broadcast and consensus that work correctly only in executions that have this structure. They point out that having a quorum system that only satisfies the consistency and availability properties can lead to liveness issues for some processes. In a slightly different model, Li et al. [17] prove that the traditional properties of quorum consistency and availability are not sufficient to solve reliable broadcast and consensus. They introduce a new property, called strong availability, very similar to the notion of a guild of Alpos et al. [2], and prove that this property is sufficient to solve the aforementioned problems. They do not comment on the necessity of such a property.
@@
Asymmetric consensus protocols have also transcended into the blockchain space with XRP Ledger (https://xprl.org[11] and Stellar (https://stellar.org[21, 18] as main representatives. In the XRP Ledger, each participant declares its own assumptions by listing other participating nodes from which it considers votes [11, 3]. Stellar uses a similar approach, in which each participant keeps a list of participants it considers important and waits for a suitable majority of them to agree on a transaction before considering it settled [21, 19, 18].
@@ 3.1 Model
All algorithms we propose are for the asynchronous unauthenticated setting. We will consider a system of $n$ processes $\mathcal{P}=\{p_{1},\dots,p_{n}\}$ that interact with each other by exchanging messages. A protocol for $\mathcal{P}$ consists of a collection of programs with instructions for all processes. They are presented using the event-based notation of Cachin et al. [10]. An execution starts with all processes in a special initial state; subsequently the processes repeatedly trigger events, react to events, and change their state through computation steps. A process that follows its protocol during an execution is called correct. A faulty process, also called Byzantine, may crash or deviate arbitrarily from its specification. We will assume that there is a low-level functionality for sending messages over point-to-point links between each pair of processes. This functionality is accessed through the events of sending and receiving a message. Point-to-point messages are authenticated and delivered reliably among correct processes. Here and from now on, the notation $\mathcal{A}^{*}$ for a system $\mathcal{A}\subseteq 2^{\mathcal{P}}$ , denotes the collection of all subsets of the sets in $\mathcal{A}$, that is, $\mathcal{A}^{*}=\{A^{\prime}|A^{\prime}\subseteq A,A\in\mathcal{A}\}$.
@@ 3.2 Asymmetric Trust Overview
We consider and expand the asymmetric trust model proposed by Alpos et al. [2]. We summarize here the necessary context for understanding this work; we refer the reader to their paper for a full presentation. In protocols with asymmetric trust, each participant is free to make its own individual trust assumptions about others, captured by an asymmetric quorum system. This contrasts with ordinary, symmetric and threshold quorum systems, where all participants share the same trust assumptions. Given a set of processes $\mathcal{P}$, an asymmetric fail-prone system $\mathbb{F}=[\mathcal{F}_{1},\dots,\mathcal{F}_{n}]$, where $\mathcal{F}_{i}$ represents the failure assumptions of process $p_{i}$, captures the heterogeneous model. Each $\mathcal{F}_{i}$ is a collection of subsets of $\mathcal{P}$ such that some $F\in\mathcal{F}_{i}$ with $F\subseteq\mathcal{P}$ is called a fail-prone set for $p_{i}$ and contains all processes that, according to $p_{i}$, may at most fail together in some execution [12]. We can, in turn, proceed to define asymmetric Byzantine quorum systems, denoted by $\mathbb{Q}$.
Definition 3.1.
@@
Given an asymmetric fail-prone system $\mathbb{F}$, there will exist a valid asymmetric quorum system for $\mathbb{F}$ if and only if $\mathbb{F}$ satisfies the $B^{3}$ condition [12, 2]. This property is defined as follows.
Definition 3.2 ($B^{3}$-condition).
@@
If $B^{3}(\mathbb{F})$ holds, then the canonical quorum system, defined as the complement of the asymmetric fail-prone system, is a valid asymmetric quorum system. This is similar to what happens in the symmetric case, where if the fail-prone system $\mathcal{F}$ satisfies the $Q^{3}$ property, the existence of a valid symmetric quorum system $\mathcal{Q}$ is guaranteed [14].
@@
Alpos et al. [2] show that naive processes may disrupt the safety and liveness guarantees of some protocols. In order to formalize this notion, they introduced the concept of a guild which is central for many of the algorithms they propose (e.g., reliable broadcast, binary consensus).
Definition 3.3.
@@
Alpos et al. [2] provide reliable broadcast and binary consensus implementations which function only in executions that contain a guild. In relation to this, Definition 3.4 presents the notion of a tolerated system, originally introduced by Alpos et al. [2]. It is built on top of tolerated sets, which are sets that contain all elements outside of a guild. This portrays the resilience of the trust assumptions, as even when all processes in one of the tolerated sets fail, there is still a guild in the system. This structure is used by Alpos et al. [2] and Zanolini [26] in their construction of several asymmetric protocols.
Definition 3.4 (Tolerated system [2]).
@@
Lemma 3.1, which is crucial for the construction presented in Section 4, shows that if there is an asymmetric fail-prone system $\mathbb{F}$ such that $B^{3}(\mathbb{F})$ holds, then $Q^{3}(\mathcal{T})$ also holds. This means that $\mathcal{T}$ is a valid symmetric fail-prone system [14].
Lemma 3.1 ([2]).
@@ 3.3 Quorum-based Algorithms
In the following sections, we will give results regarding the existence of algorithms to solve problems such as reliable broadcast and consensus. To do this, we must first define what we mean by an algorithm. We will consider all algorithms to be quorum-based algorithms. Definition 3.5 formalizes this notion. It is derived from the definitions used by Losa et al. [19] and Li et al. [17] , but it is adapted to fit the asymmetric model of Alpos et al. [2]. We use the term initiator of the execution to refer to the sole process that triggers the execution of an algorithm. In our work, this only applies within the context of reliable broadcast, where the initiator is the process that broadcasts its value. In consensus, there is no single initiator because all processes must propose a value.
Definition 3.5 (Quorum-based algorithm).
@@ 4 Guild Assumptions are not Asymmetric
In the asymmetric trust setting, Li et al. [17] show that a quorum system satisfying the classical properties of consistency and availability alone does not suffice to solve fundamental problems such as reliable broadcast or consensus. The heterogeneous views of the system make it necessary to assume more structure. Different approaches have reached similar conclusions about the assumptions needed to address these problems under asymmetric trust. Alpos et al. [2] require the existence of a guild, while Li et al. [17] propose that the system must satisfy the strong availability property. This is closely related to the idea of a guild, as it also requires the existence of a well-behaved group of processes that includes a quorum for every participant. In this section, we examine the implications of these assumptions,focusing on the model proposed by Alpos et al. [2].
@@
In the crash-fault model, Senn and Cachin [24] already show that given an asymmetric fail-prone system $\mathbb{F}$, it is possible to construct a symmetric fail-prone system $\mathcal{F}$ from $\mathbb{F}$ such that if all processes adopt $\mathcal{F}$ as their trust assumption, no process will be worse off. This result implies that there is no need for asymmetric trust in that setting. We extend this by showing that a similar scenario occurs in the Byzantine fault model for all executions with a guild. This effectively invalidates the advantages of asymmetric trust for the algorithms proposed by Alpos et al. [2] which rely on the existence of a guild.
We now show how to construct a symmetric fail-prone system $\mathcal{F}$ from an asymmetric fail-prone system $\mathbb{F}$, such that in all executions with a guild, $\mathcal{F}$ is as expressive as $\mathbb{F}$. We will build it based on the notion of the tolerated system, presented in Section 2 and used by Alpos et al. [2] and Zanolini [26] in their consensus protocols.
Definition 4.1 (Tolerated Symmetric Fail-Prone System).
Let $\mathbb{F}$ be an asymmetric fail-prone system that satisfies $B^{3}(\mathbb{F})$, $\mathbb{Q}$ be its associated canonical quorum system and $\mathcal{T}$ be the tolerated system derived from $\mathbb{Q}$ according to Definition 3.4. We interpret $\mathcal{T}$ as a symmetric fail-prone set and denote it by $\hat{\mathcal{F}}$, that is, $\hat{\mathcal{F}}=\mathcal{T}$, with a one-to-one correspondence between the tolerated sets and fail-prone sets.
From Lemma 3.1 we know that the tolerated symmetric fail-prone system satisfies the $Q^{3}$ property and therefore there will be at least one valid symmetric Byzantine quorum system associated to it.
Lemma 4.1.
Let $\mathbb{F}$ be an asymmetric fail-prone system that satisfies $B^{3}(\mathbb{F})$, $\mathcal{T}$ be the tolerated system derived from $\mathbb{F}$ and $\hat{\mathcal{F}}$ be the tolerated symmetric fail-prone system derived according to Definition 4.1. Given an execution with a guild $\mathcal{G}$ and faulty processes $F$, then $F\in\hat{\mathcal{F}}$.
@@
By Definition 3.4, there exists a tolerated set $T=\mathcal{P}\setminus\mathcal{G}$ such that $T\in\mathcal{T}$. Since $F\cap\mathcal{G}=\emptyset$ we know that $F\subseteq T$ and consequently there exists a fail-prone set $F^{\prime}\in\hat{\mathcal{F}}$ such that $T=F^{\prime}$. Therefore, the set of faulty processes will be considered by the fail-prone system $\hat{\mathcal{F}}$. ∎
Theorem 4.1.
@@
Theorem 4.1 shows that for executions with a guild, every process is at least as well off using the symmetric quorum system $\mathcal{F}$ instead of $\mathbb{F}$. For a wise process $p_{i}$ the situation will remain the same regardless of the quorum system chosen. With the asymmetric system $F\in\mathcal{F}^{*}_{i}$ will hold, while with the symmetric alternative $F\in\mathcal{F}^{*}$. For naive processes the situation improves, as now the faulty processes will be considered in their trust assumption. Thus, there is no reason for a process not to adopt the derived symmetric fail-prone system $\mathcal{F}$. Building $\mathcal{F}$ can be computationally expensive because it requires computing $\mathcal{T}$. However, existing asymmetric consensus protocols [2, 26] also require computing $\mathcal{T}$, so our approach is no more costly than current solutions.
@@
One limitation of this result is that it assumes knowledge of the trust assumptions for all processes ($\mathbb{F}$). Although this assumption holds in the Alpos et al. [2] model, it is not required in other asymmetric trust models [19]. We also note that knowledge of the faulty processes $F$ is not required to derive the symmetric fail-prone system presented in Theorem 4.1.

### v2

@@ 1 Introduction
Asymmetric trust models, such as those proposed by Damgård et al. [DBLP:conf/asiacrypt/DamgardDFN07], Alpos et al. [DBLP:journals/dc/AlposCTZ24], and Losa et al. [DBLP:conf/wdag/LosaGM19] allow the development of distributed protocols in which each participant can operate under its own trust settings. These algorithms are built on top of asymmetric quorum systems, where each process independently defines its own quorums. The traditional consistency and availability properties must be satisfied by these systems, but as was shown by Li et al. [DBLP:conf/wdag/LiCL23], they are not enough to solve fundamental problems like reliable broadcast and consensus. Therefore, the models introduce additional assumptions on the quorums. The protocols for reliable broadcast and consensus of Alpos et al. [DBLP:journals/dc/AlposCTZ24] require the existence of a guild, a set of processes that contains a quorum for each of its members. Li et al. [DBLP:conf/wdag/LiCL23] and Losa et al. [DBLP:conf/wdag/LosaGM19] identify similar conditions. These assumptions can be very strong and restrictive, and while they are sufficient to provide solutions to the aforementioned problems, we will show that they are not necessary. This leads to asymmetric algorithms that work under weaker assumptions, making such systems more flexible and usable.
We start by showing that current assumptions for algorithms such as reliable broadcast or consensus are so restrictive that it becomes unnecessary to use asymmetric trust models altogether. That is, given an asymmetric trust assumption that satisfies such requirements, we can derive an equivalent symmetric trust assumption such that no process will be worse off by adopting it. This result is surprising, considering that Alpos et al. [DBLP:journals/dc/AlposCTZ24] show that asymmetric trust is more expressive than symmetric trust.
@@ 2 Related Work
Threshold and symmetric quorum systems are fundamental tools in distributed computing, commonly used to establish the safety and liveness guarantees of various protocols [DBLP:journals/siamcomp/NaorW98, DBLP:journals/dc/MalkhiR98, DBLP:books/daglib/0017536, DBLP:books/daglib/0025983]. They constitute a crucial component of many systems in cloud computing [DBLP:journals/sigops/LakshmanM10, DBLP:conf/usenix/HuntKJR10] and cryptocurrencies [DBLP:conf/sosp/GiladHMVZ17, DBLP:journals/corr/abs-1807-04938, 1]. One drawback is that all participants must use the same quorum system, which restricts their freedom to choose who to trust. The asymmetric model, introduced by Damgård et al. [DBLP:conf/asiacrypt/DamgardDFN07] and further developed by Alpos et al. [DBLP:journals/dc/AlposCTZ24], represents a novel approach that generalizes the symmetric paradigm. It allows participants to make trust choices of their own, reflecting social connections or other information available to them outside the protocol. Each participant can specify its own quorum system, giving it the freedom to have its own trust assumptions. Amores-Sesar et al. [DBLP:conf/podc/Amores-SesarCVZ25] show that algorithms relying on symmetric quorums cannot be applied directly to the asymmetric model, as certain properties do not hold if the underlying quorums are asymmetric. This drives the need to adapt existing symmetric protocols and redefine protocol properties to fit the new model.
@@
Existing work [DBLP:journals/dc/AlposCTZ24] introduces asymmetric versions of relevant primitives like reliable broadcast, binary consensus, and common coin. Amores-Sesar et al. [DBLP:conf/podc/Amores-SesarCVZ25] extend this work by developing a DAG-based consensus protocol that relies on such primitives. Losa et al. [DBLP:conf/wdag/LosaGM19] present another approach to model asymmetric trust. They forego the notion of fail-prone sets and strengthen the definition of quorums, by requiring each quorum to contain a quorum for each member. Li et al. [DBLP:conf/wdag/LiCL23] expand on this model and address the topic of which quorum properties are necessary or sufficient to solve core asymmetric problems such as consensus. Sheff et al. [DBLP:conf/opodis/SheffWRM20] propose a variant of the Paxos protocol that incorporates heterogeneous trust assumptions. Senn and Cachin [DBLP:conf/papoc/SennC25] explore asymmetric trust in the crash-fault scenario and show that, in this setting, asymmetric trust reduces to symmetric trust.
@@
Asymmetric trust introduces novel challenges to the design of distributed protocols. One significant hurdle is that of processes that have wrong friends, that is, the existence of processes that are correct and behave according to the underlying algorithm, but whose trust assumptions do not capture the actual faulty processes. Alpos et al. [DBLP:journals/dc/AlposCTZ24] use the term naive to refer to such participants and point out that their presence can lead to safety and liveness issues in asymmetric protocols. To address this, they introduce the concept of a guild, a set of processes that contains a quorum for each member, and provide algorithms for reliable broadcast and consensus that work correctly only in executions that have this structure. They point out that having a quorum system that only satisfies the consistency and availability properties can lead to liveness issues for some processes. In a slightly different model, Li et al. [DBLP:conf/wdag/LiCL23] prove that the traditional properties of quorum consistency and availability are not sufficient to solve reliable broadcast and consensus. They introduce a new property, called strong availability, very similar to the notion of a guild, and prove that this property is sufficient to solve the aforementioned problems. They do not comment on the necessity of such a property.
@@
Asymmetric consensus protocols have also transcended into the blockchain space with XRP Ledger [DBLP:journals/corr/abs-1802-07242](https://xrpl.org) and Stellar [3, DBLP:conf/sosp/LokhavaLMHBGJMM19](https://stellar.org) as main representatives. In the XRP Ledger, each participant declares its own assumptions by listing other participating nodes from which it considers votes [DBLP:journals/corr/abs-1802-07242, DBLP:conf/opodis/Amores-SesarCM20]. Stellar uses a similar approach, in which each participant keeps a list of participants it considers important and waits for a suitable majority of them to agree on a transaction before considering it settled [3, DBLP:conf/wdag/LosaGM19, DBLP:conf/sosp/LokhavaLMHBGJMM19].
@@ 3.1 Model
All algorithms we propose are for the asynchronous unauthenticated setting. We will consider a system of $n$ processes $\mathcal{P}=\{p_{1},\dots,p_{n}\}$ that interact with each other by exchanging messages. A protocol for $\mathcal{P}$ consists of a collection of programs with instructions for all processes. They are presented using the event-based notation of Cachin et al. [DBLP:books/daglib/0025983]. An execution starts with all processes in a special initial state; subsequently the processes repeatedly trigger events, react to events, and change their state through computation steps. A process that follows its protocol during an execution is called correct. A faulty process, also called Byzantine, may crash or deviate arbitrarily from its specification. We will assume that there is a low-level functionality for sending messages over point-to-point links between each pair of processes. This functionality is accessed through the events of sending and receiving a message. Point-to-point messages are authenticated and delivered reliably among correct processes. Here and from now on, the notation $\mathcal{A}^{*}$ for a system $\mathcal{A}\subseteq 2^{\mathcal{P}}$ , denotes the collection of all subsets of the sets in $\mathcal{A}$, that is, $\mathcal{A}^{*}=\{A^{\prime}|A^{\prime}\subseteq A,A\in\mathcal{A}\}$.
@@ 3.2 Asymmetric Trust Overview
We consider and expand the asymmetric trust model proposed by Alpos et al. [DBLP:journals/dc/AlposCTZ24]. We summarize here the necessary context for understanding this work; we refer the reader to their paper for a full presentation. In protocols with asymmetric trust, each participant is free to make its own individual trust assumptions about others, captured by an asymmetric quorum system. This contrasts with ordinary, symmetric and threshold quorum systems, where all participants share the same trust assumptions. Given a set of processes $\mathcal{P}$, an asymmetric fail-prone system $\mathbb{F}=[\mathcal{F}_{1},\dots,\mathcal{F}_{n}]$, where $\mathcal{F}_{i}$ represents the failure assumptions of process $p_{i}$, captures the heterogeneous model. Each $\mathcal{F}_{i}$ is a collection of subsets of $\mathcal{P}$ such that some $F\in\mathcal{F}_{i}$ with $F\subseteq\mathcal{P}$ is called a fail-prone set for $p_{i}$ and contains all processes that, according to $p_{i}$, may at most fail together in some execution [DBLP:conf/asiacrypt/DamgardDFN07]. We can, in turn, proceed to define asymmetric Byzantine quorum systems, denoted by $\mathbb{Q}$.
Definition 1.
@@
Given an asymmetric fail-prone system $\mathbb{F}$, there will exist a valid asymmetric quorum system for $\mathbb{F}$ if and only if $\mathbb{F}$ satisfies the $B^{3}$ condition [DBLP:conf/asiacrypt/DamgardDFN07, DBLP:journals/dc/AlposCTZ24]. This property is defined as follows.
Definition 2 ($B^{3}$-condition).
@@
If $B^{3}(\mathbb{F})$ holds, then the canonical quorum system, defined as the complement of the asymmetric fail-prone system, is a valid asymmetric quorum system. This is similar to what happens in the symmetric case, where if the fail-prone system $\mathcal{F}$ satisfies the $Q^{3}$ property, the existence of a valid symmetric quorum system $\mathcal{Q}$ is guaranteed [DBLP:journals/joc/HirtM00].
@@
Alpos et al. [DBLP:journals/dc/AlposCTZ24] show that naive processes may disrupt the safety and liveness guarantees of some protocols. In order to formalize this notion, they introduced the concept of a guild which is central for many of the algorithms they propose (e.g., reliable broadcast, binary consensus).
Definition 3.
@@
Alpos et al. [DBLP:journals/dc/AlposCTZ24] provide reliable broadcast and binary consensus implementations which function only in executions that contain a guild. In relation to this, Definition 4 presents the notion of a tolerated system, originally introduced by Alpos et al. [DBLP:journals/dc/AlposCTZ24]. It is built on top of tolerated sets, which are sets that contain all elements outside of a guild. This portrays the resilience of the trust assumptions, as even when all processes in one of the tolerated sets fail, there is still a guild in the system. This structure is used by Alpos et al. [DBLP:journals/dc/AlposCTZ24] and Zanolini [4] in their construction of several asymmetric protocols.
Definition 4 (Tolerated system [DBLP:journals/dc/AlposCTZ24]).
@@
Lemma 1, which is crucial for the construction presented in Section 4, shows that if there is an asymmetric fail-prone system $\mathbb{F}$ such that $B^{3}(\mathbb{F})$ holds, then $Q^{3}(\mathcal{T})$ also holds. This means that $\mathcal{T}$ is a valid symmetric fail-prone system [DBLP:journals/joc/HirtM00].
Lemma 1 ([DBLP:journals/dc/AlposCTZ24]).
@@ 3.3 Quorum-based Algorithms
In the following sections, we will give results regarding the existence of algorithms to solve problems such as reliable broadcast and consensus. To do this, we must first define what we mean by an algorithm. We will consider all algorithms to be quorum-based algorithms. Definition 5 formalizes this notion. It is derived from the definitions used by Losa et al. [DBLP:conf/wdag/LosaGM19] and Li et al. [DBLP:conf/wdag/LiCL23] , but it is adapted to fit the asymmetric model of Alpos et al. [DBLP:journals/dc/AlposCTZ24]. We use the term initiator of the execution to refer to the sole process that triggers the execution of an algorithm. In our work, this only applies within the context of reliable broadcast, where the initiator is the process that broadcasts a value. In consensus, there is no single initiator because all processes propose a value.
Definition 5 (Quorum-based algorithm).
@@ 4 Guild Assumptions are not Asymmetric
In the asymmetric trust setting, Li et al. [DBLP:conf/wdag/LiCL23] show that a quorum system satisfying the classical properties of consistency and availability alone does not suffice to solve fundamental problems such as reliable broadcast or consensus. The heterogeneous views of the system make it necessary to assume more structure. Different approaches have reached similar conclusions about the assumptions needed to address these problems under asymmetric trust. Alpos et al. [DBLP:journals/dc/AlposCTZ24] require the existence of a guild, while Li et al. [DBLP:conf/wdag/LiCL23] propose that the system must satisfy the strong availability property. This is closely related to the idea of a guild, as it also requires the existence of a well-behaved group of processes that includes a quorum for every participant. In this section, we examine the implications of these assumptions, focusing on the model proposed by Alpos et al. [DBLP:journals/dc/AlposCTZ24].
@@
In the crash-fault model, Senn and Cachin [DBLP:conf/papoc/SennC25] already show that given an asymmetric fail-prone system $\mathbb{F}$, it is possible to construct a symmetric fail-prone system $\mathcal{F}$ from $\mathbb{F}$ such that if all processes adopt $\mathcal{F}$ as their trust assumption, no process will be worse off. This result implies that there is no need for asymmetric trust in that setting. We extend this by showing that a similar scenario occurs in the Byzantine fault model for all executions with a guild. This effectively invalidates the advantages of asymmetric trust for the algorithms proposed by Alpos et al. [DBLP:journals/dc/AlposCTZ24] which rely on the existence of a guild.
We now show how to construct a symmetric fail-prone system $\mathcal{F}$ from an asymmetric fail-prone system $\mathbb{F}$, such that in all executions with a guild, $\mathcal{F}$ is as expressive as $\mathbb{F}$. We will build it based on the notion of the tolerated system, presented in Section 2 and used by Alpos et al. [DBLP:journals/dc/AlposCTZ24] and Zanolini [4] in their consensus protocols.
Definition 6 (Tolerated Symmetric Fail-Prone System).
Let $\mathbb{F}$ be an asymmetric fail-prone system that satisfies $B^{3}(\mathbb{F})$, $\mathbb{Q}$ be its associated canonical quorum system and $\mathcal{T}$ be the tolerated system derived from $\mathbb{Q}$ according to Definition 4. We interpret $\mathcal{T}$ as a symmetric fail-prone set and denote it by $\hat{\mathcal{F}}$, that is, $\hat{\mathcal{F}}=\mathcal{T}$, with a one-to-one correspondence between the tolerated sets and fail-prone sets.
From Lemma 1 we know that the tolerated symmetric fail-prone system satisfies the $Q^{3}$ property and therefore there will be at least one valid symmetric Byzantine quorum system associated to it.
Lemma 2.
Let $\mathbb{F}$ be an asymmetric fail-prone system that satisfies $B^{3}(\mathbb{F})$, $\mathcal{T}$ be the tolerated system derived from $\mathbb{F}$ and $\hat{\mathcal{F}}$ be the tolerated symmetric fail-prone system derived according to Definition 6. Given an execution with a guild $\mathcal{G}$ and faulty processes $F$, then $F\in\hat{\mathcal{F}}$.
@@
By Definition 4, there exists a tolerated set $T=\mathcal{P}\setminus\mathcal{G}$ such that $T\in\mathcal{T}$. Since $F\cap\mathcal{G}=\emptyset$ we know that $F\subseteq T$ and consequently there exists a fail-prone set $F^{\prime}\in\hat{\mathcal{F}}$ such that $T=F^{\prime}$. Therefore, the set of faulty processes will be considered by the fail-prone system $\hat{\mathcal{F}}$. ∎
Theorem 1.
@@
Theorem 1 shows that for executions with a guild, every process is at least as well off using the symmetric quorum system $\mathcal{F}$ instead of $\mathbb{F}$. For a wise process $p_{i}$ the situation will remain the same regardless of the quorum system chosen. With the asymmetric system $F\in\mathcal{F}^{*}_{i}$ will hold, while with the symmetric alternative $F\in\mathcal{F}^{*}$. For naive processes the situation improves, as now the faulty processes will be considered in their trust assumption. Thus, there is no reason for a process not to adopt the derived symmetric fail-prone system $\mathcal{F}$. Building $\mathcal{F}$ can be computationally expensive because it requires computing $\mathcal{T}$. However, existing asymmetric consensus protocols [DBLP:journals/dc/AlposCTZ24, 4] also require computing $\mathcal{T}$, so our approach is no more costly than current solutions.
@@
One limitation of this result is that it assumes knowledge of the trust assumptions for all processes ($\mathbb{F}$). Although this assumption holds in the Damgård et al. [DBLP:conf/asiacrypt/DamgardDFN07] and Alpos et al. [DBLP:journals/dc/AlposCTZ24] models, it is not required in other asymmetric trust models [DBLP:conf/wdag/LosaGM19]. We also note that knowledge of the faulty processes $F$ is not required to derive the symmetric fail-prone system presented in Theorem 1.